OpenStudy (anonymous):

what is the derivative of f(x)= -2|x|^1/2

OpenStudy (anonymous):

use the chain rule: d/dx (-2 abs(x)^(1/2)) =-2 d/dx(abs(x)^(1/2)) =-2 * 1/2 * abs(x)^(-1/2) * d/dx (abs(x)) =-abs'(x)/sqrt(abs(x)) abs'(x)=x/abs(x); =-x/(|x| sqrt(|x|))

OpenStudy (anonymous):

this assumes the domain of f is the reals; f(x) is not a holomorphic function

OpenStudy (anonymous):

so if we have 2|x|^1/2 would it look like: =2 d/dx(abs(x)^(1/2)) =2 * 1/2 * abs(x)^(-1/2) * d/dx (abs(x)) =abs'(x)/sqrt(abs(x)) =abs'(x)=x/abs(x); =x/(|x| sqrt(|x|)) or did i mess something up? @eashy

OpenStudy (anonymous):

looks good :)

OpenStudy (anonymous):

thanks